# Learning to crawl: benefits and limits of centralized vs distributed control

### v1

@@ Abstract
We present a model of a crawler consisting of several suckers distributed along a straight line and connected by springs. Both proprioception and control are binary: the system responds to the elongation vs compression of its springs and the suckers can either adhere or remain idle. A central pattern generator delivers a traveling wave of compression, and the crawler is tasked to learn how to control adhesion to effectively ride the wave in order to crawl. We ask what are the benefits and limitations of distributed vs centralized learning architectures. Using tabular Q-learning we demonstrate that crawling can be learned by trial and error in a purely distributed setting where each sucker learns independently how to control adhesion, with no exchange of information among them. We find that centralizing control of all suckers enhances speed and robustness to failure by orchestrating smoother collective dynamics and partly overcoming the limitations of rudimentary proprioception and control. However, since the computational cost scales exponentially with the number of suckers, centralized control quickly becomes untreatable with the size of the crawler. We then show that intermediate levels of centralization, where multiple control centers coordinate subsets of suckers, can negotiate fast and robust crawling while avoiding excessive computational burden. Our model can be further generalized to explore the trade-offs between crawling speed, robustness to failure, computational cost and information exchange that shape the emergence of biological solutions for crawling and could inspire the design of robotic crawlers.
@@ I Introduction
These nervous systems showcase different architectures reflecting varying degrees of distribution vs centralization. In cnidarians, the nervous system consists only of diffuse nerve nets and is the most primitive of all animals’. Despite the lack of centralization, cnidaria are capable of complex behaviors. For example, Hydra vulgaris achieves highly non-trivial coordination of its body parts to perform somersaults [1]. Furthermore, evidence of associative learning in the form of operant conditioning was shown for the box jellyfish T. Cystophora, induced by either visual or mechanical stimuli [2].
@@
Echinoderms, such as starfish, also have a simple radial nervous system consisting of a nerve net of neurons. But while they have no brain, their nervous system is more structured than in the cnidarians, as the nerves connect to a central nerve ring [3]. Starfishes crawl efficiently using hundreds of independent tubular feet which generate robust locomotion without central control. The nerve ring establishes the direction of motion by integrating sensory information detected by the arms [4, 5]. Given the direction of motion, coordination of the tube feet can emerge purely from a passive mechanical coupling, as showed with a minimal model where no information propagates along the nerve net [6, 7].
@@
Cephalopods, such as octopus, squid, and cuttle fish, have evolved a complex nervous system whose size is comparable to that of vertebrates and endows them with impressive cognitive capabilities. They do have a central brain; however most of their nervous system is distributed throughout their arms which are covered in hundreds of suckers that can sense the environment as well as adhere to substrates. Like all soft-bodied animals, these organisms have to coordinate many degrees of freedom to achieve locomotion. For example, octopuses translocate on surfaces by leveraging their taste-by-touch sense, whereby suckers adhere to surfaces and sense them chemically upon contact [8]. The sensory information acquired by the periphery guides suckers’ adhesion, yielding displacement of the organism along the surface. However, it is not completely understood to what extent locomotion is controlled by the central nervous system vs by the ganglia distributed along the arms [9, 10]. Intriguingly, coordination may or may not depend on precise proprioceptive feedback, since the molecular mechanisms for proprioception in cephalopods are unknown [11, 12].
@@
We draw inspiration from these natural crawlers and focus on a toy model to ask what are the pros and cons of centralization vs distribution. The crawler is modeled as a collection of suckers distributed along a straight line and connected by springs, and endowed with a rudimentary form of proprioception (extension vs elongation of the springs) and a rudimentary form of control of the suckers (adhere vs not adhere). A central pattern generator (CPG) prompts a traveling wave of compression of the springs. If suckers adhere at random times, the crawler center of mass cannot move on average: can we learn a pattern of adhesion that enables net translocation?
@@
To answer this question we develop a simulator of the crawler and let it learn to translocate by trial and error using a reinforcement learning algorithm called Q-learning. We compare the learnability, performance and robustness of centralized vs distributed learning architectures. Despite the rudimentary proprioception and control, we find that purely distributed architectures, where each sucker is an independent agent, learn to translocate. Centralized control, where all suckers are simultaneously controlled by a single agent, achieves faster crawling and increased robustness to failure, partially overcoming the crawler’s poor sensorimotor ability. However, the computational cost of the central control scales exponentially with the number of suckers and springs. We then introduce a partially centralized architecture and find that it achieves nearly optimal performance and robustness with a limited increase in computational burden. These results complement previous work that focus on either a fully centralized [13] or a fully distributed [14] learning architecture as well as works where crawling emerges from prescribed control patterns, rather than learning (see e.g. [15, 6])
@@
Our toy model of crawling shows that the sole existence of a common reward –here the speed of the center of mass– is sufficient for independent suckers to coordinate. However, when information and control are shared, crawling is faster and more reliable. We hope that our work will serve as a benchmark to explore the optimal level of centralization in more realistic scenarios for natural and artificial crawling, where constraints on efficiency, robustness and computational costs will all concur to define optimality.
@@ IV Discussion and perspectives
In contrast, previous biologically inspired work focused on either a fully centralized [13] or fully distributed (hive) control [14]. In [13], a biologically plausible model of crawling is proposed which also uses rudimentary forms of proprioception and control: input is limited to the identity of the most compressed body segment and actions correspond to the designation of the sole active neuron in the neural network controlling muscular contraction. The authors show with a fully centralized Q–learning architecture that the central pattern generator can be learned, rather than imposed as we do here.
In [14], swimming of an agent is achieved through training of a genetic algorithm with a fully distributed architecture, where individual actuators are all forced to learn the same policy – which we have called hive update here. In their work, the actuators are not suckers but rather beads which control the restoring force between their neighbors. The hive update was found to generalize well to a different number of beads, with no need for re-training. This is consistent with our finding, where we also observe that the hive update almost invariably learns the same policy, independent on the number of suckers composing the crawler. However, while we focus on binary sensing and control, Ref. [14] considers actuators with full proprioception, including the position and speeds of their neighbors. An advantage of considering binary proprioception is its simple implementation and interpretation. For example, our hive update can be understood in simple terms: suckers learn to only adhere when they are pushed and pulled backward from both sides.
Here we have focused on the learning process that enables suckers to achieve crawling by combining mechanical contractions and interaction with the substrate. In our model, thrust is achieved by letting suckers control patterns of adhesion to the solid substrate, whereas contraction is prescribed by the dynamics. Learning is also the focus in refs. [13, 14], with a switch in the controls: the actuators learn patterns of contraction, and interaction with their solid and fluid environment respectively is embedded in the dynamics. In contrast, thrust is achieved by fully prescribing –rather than learning– actions in a bio-inspired model of sea star locomotion [7, 6]. Synchronization here emerges through the dynamical interaction of the substrate with the tube feet, which are all physically connected to the body, with a mechanism reminiscent of the Kuramoto model of synchronization [27, 28].

### v2

@@ Abstract
We present a model of a crawler consisting of several suction units distributed along a straight line and connected by springs. The suction units are rudimentary proprioceptors-actuators, which sense binary states of compression vs elongation of the springs, and can either adhere or remain idle. Muscular contraction is not controlled by the crawler, but follows an endogenous, stereotyped wave. The crawler is tasked to learn patterns of adhesion that generate thrust in response to the wave of contraction. Using tabular Q-learning we demonstrate that crawling can be learned by trial and error and we ask what are the benefits and limitations of distributed vs centralized learning architectures. We find that by centralizing proprioceptive feedback and control, the crawler leverages long range correlations in the dynamics and ride the endogenous wave smoothly. The ensuing benefits are measured in terms of both speed and robustness to failure, although they come at increased computational cost. At the opposite extreme, purely distributed feedback and control only leverages local information and yield a jerkier and slower crawling, although computationally cheap. Intermediate levels of centralization can negotiate fast and robust crawling while avoiding excessive computational burden, demonstrating the computational benefits of a hierarchical organization of crawling. Our model unveils the trade-offs between crawling speed, robustness to failure, computational cost and information exchange that may shape biological solutions for crawling and could inspire the design of robotic crawlers.
@@ I Introduction
These nervous systems showcase different architectures reflecting varying degrees of distribution vs centralization. In cnidarians, the nervous system consists only of diffuse nerve nets and is the most primitive of all animals’. Despite the lack of centralization, cnidaria are capable of complex behaviors. For example, Hydra vulgaris achieves highly non-trivial coordination of its body parts to perform somersaults [1]. Furthermore, associative learning in the form of operant conditioning can be induced by either visual or mechanical stimuli for the box jellyfish T. Cystophora [2].
@@
Echinoderms, such as starfish, also have a simple radial nervous system consisting of a nerve net of neurons. But while they have no brain, their nervous system is more structured than in the cnidarians, as the nerves connect to a central nerve ring [3]. Starfishes crawl efficiently using hundreds of independent tubular feet which generate robust locomotion without central control. The nerve ring establishes the direction of motion by integrating sensory information detected by the arms [4, 5]. Given the direction of motion, coordination of the tube feet can emerge purely from a passive mechanical coupling, as showed with a minimal model where behavior of the tube feet is prescribed and no information propagates along the nerve net [6, 7].
@@
Cephalopods, such as octopus, squid and cuttlefish, have evolved a complex nervous system whose size is comparable to that of vertebrates and endows them with impressive cognitive capabilities. They do have a central brain; however most of their nervous system is distributed throughout their arms which are covered in hundreds of suckers that can sense the environment as well as adhere to substrates. Like all soft-bodied animals, these organisms have to coordinate many degrees of freedom to achieve locomotion. For example, octopuses translocate on surfaces by leveraging their taste-by-touch sense, whereby suckers adhere to surfaces and sense them chemically upon contact [8]. The sensory information acquired by the periphery guides suckers’ adhesion, yielding displacement of the organism along the surface. However, to what extent locomotion is controlled by the central nervous system vs by the ganglia distributed along the arms is not known [9, 10]. Intriguingly, coordination may or may not depend on precise proprioceptive feedback, since the molecular mechanisms for proprioception in cephalopods are unknown [11, 12].
@@
We draw inspiration from these natural crawlers and focus on a toy model to ask what are the pros and cons of centralization vs distribution. The crawler is modeled as a collection of suckers distributed along a straight line and connected by springs, and endowed with a rudimentary form of proprioception (extension vs elongation of the springs) and a rudimentary form of control of the suckers (adhere vs not adhere). Contraction of the springs is not controlled by the crawler but rather follows a stereotyped wave established by a central pattern generator (CPG). If suckers were to adhere at random times, the crawler center of mass would be unable to move on average: is it possible to learn a pattern of adhesion that enables net translocation?
@@
To answer this question we develop a simulator of the crawler and let it learn to translocate by trial and error using the Q-learning Reinforcement Learning algorithm. We compare the learnability, performance and robustness of centralized vs distributed learning architectures. Despite the rudimentary proprioception and control, we find that the sole existence of a common reward –here the speed of the center of mass– is sufficient for independent suckers to learn crawling. When all suckers are simultaneously controlled by a single agent, crawling is both faster and more robust to failure, partially overcoming the crawler’s poor sensorimotor ability. However, the computational cost of the central control scales exponentially with the number of suckers and springs. We then introduce a shallow hierarchical architecture where control is centralized in few control centers. We find that this partial centralization achieves nearly optimal performance and robustness with a limited increase in computational burden. These results complement previous work that focus on controlling contraction rather than adhesion with either purely centralized [13] or purely distributed [14] architectures as well as works where crawling emerges from prescribed control patterns, rather than learning (see e.g. [15, 6])
@@
We hope that our work will serve as a benchmark to explore the optimal level of centralization in more realistic scenarios for natural and artificial crawling, where constraints on efficiency, robustness and computational costs will all concur to define optimality.
@@ IV Discussion and perspectives
In contrast, previous biologically inspired work focused on either a fully centralized [13] or fully distributed (hive) control [14]. In [13], a biologically plausible model of crawling is proposed which also uses rudimentary forms of proprioception and control: input is limited to the identity of the most compressed body segment and actions correspond the designation of the sole active neuron in the neural network controlling muscular contraction, while adhesion is not controlled. The authors show with a fully centralized Q–learning architecture that the central pattern generator can be learned, rather than imposed as we do here.
In [14], swimming of an agent is achieved through training of a genetic algorithm with a fully distributed architecture, where individual actuators are all forced to learn the same policy – which we have called hive update here. Again, the actuators are not suckers but rather beads which control the restoring force between their neighbors. The hive update was found to generalize well to a different number of beads, with no need for re-training. This is consistent with our finding, where we also observe that the hive update almost invariably learns the same policy, independent on the number of suckers composing the crawler. However, while we focus on binary sensing and control, Ref. [14] considers actuators with full proprioception, including the position and speeds of their neighbors. An advantage of considering binary proprioception is its simple implementation and interpretation. For example, our hive update can be understood in simple terms: suckers learn to only adhere when they are pushed and pulled backward from both sides.
Here we have focused on the learning process that enables suckers to achieve crawling by combining mechanical contractions and interaction with the substrate. In our model, thrust is achieved by letting suckers control patterns of adhesion to the solid substrate, whereas contraction results from the stereotypic wave of compression and the dynamics. In refs. [13, 14] switch the controls: the actuators learn patterns of contraction, and interaction with their solid and fluid environment respectively is embedded in the dynamics. Finally, in bio-inspired models of sea star locomotion [7, 6], thrust is achieved by fully prescribing –rather than learning– actions. Synchronization here emerges through the dynamical interaction of the substrate with the tube feet, which are physically connected to the body, with a mechanism reminiscent of the Kuramoto model of synchronization [27, 28].